OpenStudy (anonymous):

HELP ME PLEASE!!!!! Solve, using the substitution method. y – 3x = 7 21 + 9x = 3y A. There is no solution. B. There are an infinite number of solutions. C. The solution is (1, 10) D. The solution is (28, 0)

jimthompson5910 (jim_thompson5910):

hint: y – 3x = 7 y – 3x+3x = 7+3x y + 0x = 3x + 7 y + 0 = 3x + 7 y = 3x + 7

OpenStudy (anonymous):

Wait I Dont get the second part

OpenStudy (anonymous):

21 + 9x = 3(3x+7)

jimthompson5910 (jim_thompson5910):

do you see how I went from y – 3x = 7 to y = 3x + 7 ?

jimthompson5910 (jim_thompson5910):

3(3x+7) distributes to 3*3x+3*7 and that turns into 9x + 21

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

21 + 9x = 3(3x+7) turns into 21 + 9x = 9x + 21

OpenStudy (anonymous):

but what do you do when its 9x + 21=9x + 21

OpenStudy (anonymous):

whats the soulution thats what im stuck on

jimthompson5910 (jim_thompson5910):

what happens when you subtract 9x from both sides

OpenStudy (anonymous):

21=21

jimthompson5910 (jim_thompson5910):

is that always true? sometimes true? or always false?

OpenStudy (anonymous):

idk

jimthompson5910 (jim_thompson5910):

is 21 equal to itself?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so this never changes, so it's always true

OpenStudy (anonymous):

so is it a b c or d

jimthompson5910 (jim_thompson5910):

because it's ALWAYS true, this means that x and y have no effect at all

jimthompson5910 (jim_thompson5910):

so x can be any real number and the system would have a solution

OpenStudy (anonymous):

so it b

OpenStudy (anonymous):

There are an infinite number of solutions

OpenStudy (anonymous):

Am i right?

jimthompson5910 (jim_thompson5910):

yes you are

OpenStudy (anonymous):

ok thanks alot

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

GUYS @jim_thompson5910 IS THE BEST!!!!!!!!!!